### v1

@@ 4 Results and Discussion
We report the performance of pretrained models, using raw images (No Seg), images with the segmentation masks applied (Seg), and the difference in performance attributable to using the segmented images ($\Delta$ Seg). On the French Census dataset, Table 1 shows that in almost every case the segmented images significantly improve performance. The exception to this is CLIP-ViT-L/14-336, which has minor changes in performance in both directions.
@@
CLIP-ViT-B/32 0.691 0.706 +0.015 0.757 0.808 +0.051 0.860 0.900 +0.040 0.959 0.973 +0.014
CLIP-ViT-L/14-336 0.833 0.804 −0.030 0.875 0.864 −0.012 0.928 0.932 +0.004 0.964 0.976 +0.012
DiT-Base 0.398 0.812 +0.413 0.565 0.855 +0.290 0.755 0.923 +0.169 0.958 0.968 +0.010
DiT-Large 0.453 0.795 +0.342 0.602 0.863 +0.261 0.773 0.931 +0.158 0.963 0.980 +0.017
ResNet18 0.722 0.799 +0.077 0.816 0.869 +0.053 0.917 0.949 +0.032 0.981 0.978 −0.003
ResNet50 0.765 0.844 +0.079 0.849 0.896 +0.047 0.948 0.953 +0.005 0.851 0.981 +0.131
@@
It is important to note that the models have different embedding dimension sizes for the linear probe, so linear probe accuracy may not be appropriate to compare across models. For instance, the poor linear probe performance of ResNet50 on the French Census dataset may be attributed to its larger embedding dimension of 2048 compared to 512 in ResNet18, 768 in MAE and DiT, and 512 for CLIP. This, however, does not detract from the $\Delta$ we observe when using the segmented images. Notably, because we have reduced the number of dimensions using UMAP before computing the other metrics, the comparison across models on K-means and KNN metrics is still valid.
@@ 4.1 Self-supervision with MAEs
Raw Image
Segmented Image
Figure 4: The segmented image shows some parts of column headers incorrectly masked out, which may contribute to diminished performance.
Table 3 shows the results on the French Census dataset. The findings indicate that MAE models trained on each dataset perform better than the pretrained ones we evaluated. This advantage is likely due, at least in part, to the increased resolution, a benefit that will persist until pretrained transformer models adopt higher resolution inputs. Segmentation does not appear to help the French Census and sometimes harms performance, suggesting that better segmentation may improve results. Figure 4 illustrates an example of imperfect segmentation that may harm performance.
@@
Segmentation 0.902 0.944 0.980 0.990
Segmentation + Translation 0.824 0.913 0.966 0.985

### v2

@@ 4 Results and Discussion
We report the performance of pretrained models, using raw images (No Seg), images with the segmentation masks applied (Seg), and the difference in performance attributable to using the segmented images ($\Delta$ Seg). On the French Census dataset, Table 1 shows that in almost every case the segmented images improve performance, often significantly. The exception to this is CLIP-ViT-L/14-336, which has minor changes in performance in both directions.
@@
CLIP-ViT-B/32 0.694 0.704 +0.011 0.757 0.806 +0.048 0.860 0.903 +0.043 0.959 0.973 +0.014
CLIP-ViT-L/14-336 0.833 0.809 −0.024 0.875 0.868 −0.008 0.928 0.938 +0.010 0.964 0.976 +0.012
DiT-Base 0.397 0.826 +0.429 0.564 0.861 +0.297 0.755 0.927 +0.172 0.958 0.968 +0.010
DiT-Large 0.444 0.794 +0.350 0.602 0.861 +0.259 0.773 0.929 +0.156 0.963 0.980 +0.017
ResNet18 0.719 0.794 +0.075 0.816 0.868 +0.052 0.917 0.949 +0.032 0.981 0.978 −0.003
ResNet50 0.772 0.864 +0.092 0.853 0.906 +0.052 0.948 0.955 +0.008 0.980 0.981 +0.002
@@
It is important to note that the models have different embedding dimension sizes for the linear probe, so linear probe accuracy may not be appropriate for comparison across models. The embedding dimensions are 2048 for ResNet50, 512 for ResNet18, 768 for MAE and DiT, and 512 for CLIP. This, however, does not detract from the $\Delta$ we observe when using the segmented images. Notably, because we have reduced the number of dimensions using UMAP before computing the other metrics, the comparison across models on K-means and KNN metrics is still valid.
@@ 4.1 Self-supervision with MAEs
Table 3 shows the results on the French Census dataset. The findings indicate that MAE models trained on each dataset perform better than the pretrained ones we evaluated. This advantage is likely due, at least in part, to the increased resolution, a benefit that will persist until pretrained transformer models adopt higher-resolution inputs. All models tested show strong KNN and linear probe accuracy, though clustering seems to improve when using segmented images and translation as an augmentation.
@@
Segmentation 0.920 0.953 0.980 0.990
Segmentation + Translation 0.939 0.963 0.989 0.997